# Recursive Gaussian Process State Space Model

### v2

@@ 5.2 Computational Complexity Analysis
Total complexity $\mathcal{O}\left((d_{x}+d_{f}M)^{3}\right)$ $\mathcal{O}\left((d_{x}+d_{f}M)^{3}\right)$
@@ 6 Experiment Result
This section demonstrates the proposed method through two synthetic simulations and a real-world dataset. Section 6.1 presents a comparative experiment on a synthetic NASCAR® dataset. Section 6.2 showcases the capability of RGPSSM for hyperparameter and inducing points adaptation in a synthetic wing rock dynamics learning task. Section 6.3 evaluates the learning performance of RGPSSM on various real-world datasets.
@@ 6.1 Synthetic NASCAR® Dynamics Learning
We evaluate the online learning performance of RGPSSM by comparing it with two representative algorithms: SVMC [39], which represents PF-based methods, and OEnVI [35], which exemplifies SVI-based methods. Since the pre-parameterized AKF-based method is a special case of the proposed method, it is not included in the comparison. In the experiment, all methods are tested using synthetic NASCAR® data [55], which involves a two-dimensional state with dynamics following a recurrent switching linear dynamical system. The measurement model is given by $\bm{y}_{t}=\bm{C}\bm{x}_{t}+\bm{w}_{t},\,\bm{w}_{t}\sim N(\bm{0},0.1^{2}\bm{I}_{4})$, where $d_{y}=4$ and $\bm{C}$ is a 4-by-2 matrix with random elements. The three methods are trained with 500 measurements and tested by predicting the subsequent 500 steps. Note that, unlike the experiments in [39] and [35], we use fewer dimensions and a smaller number of measurements for training, providing a more rigorous evaluation. For a fair comparison, the maximum number of inducing points across all methods is limited to 20, and the GP hyperparameters are set to the same values and are not adjusted online. The implementation and other parameter settings for the SVMC and OEnVI algorithms are based on the code provided online [^9] https://github.com/catniplab/svmc  [^10] https://github.com/zhidilin/gpssmProj.
@@ 6.2 Synthetic Wing Rock Dynamics Learning
To further evaluate RGPSSM’s ability to online adapt GP hyperparameters and inducing points, it is tested on the synthetic wing rock dynamics learning task. In aerospace engineering, wing rock dynamics can cause roll oscillations in slender delta-winged aircraft, posing a significant threat to flight safety. To address this issue, most approaches rely on online learning of wing rock dynamics [56, 49]. The associated dynamical system can be expressed by the following continuous-time SSM:
@@
where $\theta,p\in\mathbb{R}$ denote the roll angle and roll rate, respectively. As illustrated in (35), the dynamics model consists of a known part $L_{\delta_{a}}\delta_{a}$, where $\delta_{a}$ is the aileron control input and $L_{\delta_{a}}=3$, and an unknown uncertainty model $\Delta(\theta,p)$. The uncertainty model used in the simulation is taken from [56], namely:
@@
where $W_{0}=0.8,W_{1}=0.2314,W_{2}=0.6918,W_{3}=-0.6245,W_{4}=0.0095,W_{5}=0.0214$. To further demonstrate the learning capacity of the proposed method under sparse sensor conditions, the measurement is limited to the roll angle $\theta$, which is corrupted by Gaussian white noise with a standard deviation of 0.2 degrees.
@@
Firstly, we conduct a Monte Carlo test of the proposed learning algorithm with different initial values of GP hyperparameters. The simulation results are depicted in Fig. 3 and 4. As shown in Fig. 3, all three GP hyperparameters gradually converge near the reference values, which are obtained through offline GP training using data pairs $(\theta,p,\Delta)$. To evaluate the improvement resulting from hyperparameter adaptation, we calculate the prediction RMSE over the last quarter segment, which is 29.3% lower with hyperparameter adaptation than without. Furthermore, to more intuitively illustrate the benefits of hyperparameter adaptation, we present the state profile from one Monte Carlo simulation case, as shown in Fig. 4. In the figure, the blue solid lines depict the true values of state $\theta(t),p(t)$ and uncertainty $\Delta(t)$; the orange dashed lines represent the filtered state and uncertainty predictions obtained without hyperparameter adaptation; and the green dashed lines show the predictions with hyperparameter adaptation. It is evident that hyperparameter adaptation improves both the mean and variance predictions of the GP, thereby enhancing the filtering accuracy of the state, especially the roll rate $p$. Therefore, the proposed method can not only learn the uncertain dynamics but also improve the state estimation accuracy.
@@ 6.3 Real-World System Identification Task
Actuator 0.691 (0.148) 0.815 (0.012) 0.657 (0.095) 1.048 (0.142)
Ball Beam 0.074 (0.010) 0.065 (0.005) 0.055 (0.002) 0.046 (0.002)
Drive 0.647 (0.057) 0.735 (0.005) 0.703 (0.050) 0.863 (0.083)
Dryer 0.174 (0.013) 0.667 (0.266) 0.125 (0.017) 0.105 (0.004)
Gas Furnace 1.503 (0.196) 2.052 (0.163) 1.388 (0.123) 1.373 (0.061)

### v3

@@ 5.2 Computational Complexity Analysis
Total complexity $\mathcal{O}\left((d_{x}+d_{f}M)^{3}\right)$ - $\mathcal{O}\left((d_{x}+d_{f}M)^{3}\right)$ -
@@ 6 Experiment Result
This section demonstrates the proposed method using two synthetic simulations and a real-world dataset. Section 6.1 presents a comparative experiment on a synthetic NASCAR® dataset, which serves as a benchmark for online GPSSM methods. Section 6.2 showcases the capability of RGPSSM for hyperparameter and inducing point adaptation in a synthetic wing rock dynamics learning task. Section 6.3 evaluates the learning performance of RGPSSM on various real-world datasets.
@@ 6.1 Synthetic NASCAR® Dynamics Learning
To benchmark the online learning performance of RGPSSM, we compare it with two state-of-the-art algorithms: SVMC [44], which represents PF-based methods, and OEnVI [39], which exemplifies SVI-based methods. Since the pre-parameterized AKF-based method is a special case of the proposed method, it is not included in the comparison. In the experiment, all methods are tested using synthetic NASCAR® data [60], which involves a two-dimensional state with dynamics following a recurrent switching linear dynamical system. The measurement model is given by $\bm{y}_{t}=\bm{C}\bm{x}_{t}+\bm{w}_{t},\,\bm{w}_{t}\sim N(\bm{0},0.1^{2}\bm{I}_{4})$, where $d_{y}=4$ and $\bm{C}$ is a 4-by-2 matrix with random elements. The three methods are trained with 500 measurements and tested by predicting the subsequent 500 steps. Note that, unlike the experiments in [44] and [39], we use fewer dimensions and a smaller number of measurements for training, providing a more rigorous evaluation. For a fair comparison, the maximum number of inducing points across all methods is limited to 20, and the GP hyperparameters are set to the same values and are not adjusted online. The implementation and other parameter settings for the SVMC and OEnVI algorithms are based on the code provided online [^9] https://github.com/catniplab/svmc  [^10] https://github.com/zhidilin/gpssmProj.
@@ 6.2 Synthetic Wing Rock Dynamics Learning
To further evaluate RGPSSM’s ability to online adapt GP hyperparameters and inducing points, it is tested on the synthetic wing rock dynamics learning task. In aerospace engineering, wing rock dynamics can cause roll oscillations in slender delta-winged aircraft, posing a significant threat to flight safety. To address this issue, most approaches rely on online learning of wing rock dynamics [6, 7, 8, 61]. The associated dynamical system can be expressed by the following continuous-time SSM:
@@
where $\theta\in\mathbb{R}$ ($\mathrm{d}\mathrm{e}\mathrm{g}$) and $p\in\mathbb{R}$ ($\mathrm{d}\mathrm{e}\mathrm{g}\mathrm{/}\mathrm{s}$) denote the roll angle and roll rate, respectively. As illustrated in (35), the dynamics model consists of a known part $L_{\delta_{a}}\delta_{a}$, where $\delta_{a}$ is the aileron control input and $L_{\delta_{a}}=3~${\mathrm{s}}^{-2}$\$, and an unknown uncertainty model $\Delta(\theta,p)$. The uncertainty model used in the simulation is taken from [61], namely:
@@
where $W_{0}=0.8\,$\mathrm{d}\mathrm{e}\mathrm{g}\cdot\mathrm{s}^{-2}$\$, $W_{1}=0.2314\,$\mathrm{s}^{-2}$\$, $W_{2}=0.6918\,$\mathrm{s}^{-1}$\$, $W_{3}=-0.6245\,$\mathrm{d}\mathrm{e}\mathrm{g}^{-1}\mathrm{s}^{-1}$\$, $W_{4}=0.0095\,$\mathrm{d}\mathrm{e}\mathrm{g}^{-1}$\$, $W_{5}=0.0214\,$\mathrm{d}\mathrm{e}\mathrm{g}^{-2}\,\mathrm{s}^{-2}$\$. To further demonstrate the learning capacity of the proposed method under sparse sensor conditions, the measurement is limited to the roll angle $\theta$, which is corrupted by Gaussian white noise with a standard deviation of 0.2 degrees.
@@
Firstly, we conduct a Monte Carlo test of the proposed learning algorithm with different initial values of GP hyperparameters. The simulation results are depicted in Fig. 3 and 4. As shown in Fig. 3, all three GP hyperparameters gradually converge near the reference values, which are obtained through offline GP training using data pairs $(\theta,p,\Delta)$. This indicates that the GP hyperparameters do not need to be precisely selected at the stage of algorithm initialization, and the online optimization can effectively adjust them. To evaluate the improvement resulting from hyperparameter adaptation, we calculate the prediction RMSE over the last quarter segment, which is 29.3% lower with hyperparameter adaptation than without. Furthermore, to more intuitively illustrate the benefits of hyperparameter adaptation, we present the state profile from one Monte Carlo simulation case, as shown in Fig. 4. In the figure, the blue solid lines depict the true values of state $\theta(t),p(t)$ and uncertainty $\Delta(t)$; the orange dashed lines represent the filtered state and uncertainty predictions obtained without hyperparameter adaptation; and the green dashed lines show the predictions with hyperparameter adaptation. It is evident that hyperparameter adaptation improves both the mean and variance predictions of the GP, thereby enhancing the filtering accuracy of the state, especially the roll rate $p$. Therefore, the proposed method can not only learn the uncertain dynamics but also improve the state estimation accuracy.
@@ 6.3 Real-World System Identification Task
Actuator
0.691
0.815
0.657
1.048
Ball Beam
0.074
0.065
0.055
0.046
Drive
0.647
0.735
0.703
0.863
Dryer
0.174
0.667
0.125
0.105
Gas Furnace
1.503
2.052
1.388
1.373